Direct Materials, Direct Labor, and Reports budgeted and actual costs for variable and fixed factory overhead...

  1. Direct Materials, Direct Labor, and Reports budgeted and actual costs for variable and fixed factory overhead along with the related controllable and volume variances.Factory Overhead Cost Variance Analysis

    Mackinaw Inc. processes a base chemical into plastic. A detailed estimate of what a product should cost.Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 70,000 units of product were as follows:

    Standard Costs Actual Costs
    Direct materials 189,000 lbs. at $6.00 187,100 lbs. at $5.90
    Direct labor 17,500 hrs. at $16.50 17,900 hrs. at $16.90
    Factory overhead Rates per direct labor hr.,
    based on 100% of normal
    capacity of 18,260 direct
    labor hrs.:
    Variable cost, $3.40 $58,910 variable cost
    Fixed cost, $5.40 $98,604 fixed cost

    Each unit requires 0.25 hour of direct labor.

    Required:

    a. Determine the direct materials Price variance is the difference between the actual and standard prices, multiplied by the actual quantity.price variance, direct materials The cost associated with the difference between the standard quantity and the actual quantity of direct materials used in producing a commodity.quantity variance, and total direct materials The difference between actual cost and standard cost at actual volumes.cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Materials Price Variance $ Favorable
    Direct Materials Quantity Variance $ Favorable
    Total Direct Materials Cost Variance $ Favorable

    b. Determine the direct labor The cost associated with the difference between the standard rate and the actual rate paid for direct labor used in producing a commodity.rate variance, direct labor The cost associated with the difference between standard and actual hours of direct labor spent for producing a commodity.time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Labor Rate Variance $ Unfavorable
    Direct Labor Time Variance $ Unfavorable
    Total Direct Labor Cost Variance $ Unfavorable

    c. Determine the variable factory overhead The difference between the actual variable overhead costs and the budgeted variable overhead for actual production.controllable variance, fixed factory overhead The difference between the budgeted fixed overhead at 100% of normal capacity and the standard fixed overhead for the actual production achieved during the period.volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Variable factory overhead controllable variance $ Favorable
    Fixed factory overhead volume variance $ Unfavorable
    Total factory overhead cost variance $ Unfavorable
Answer #1
Direct Materials Price Variance [(Actual price-Standard Price)*Actual Quantity]
=(5.90-6)*187100    (18,710) F
Direct Materials Quantity Variance [(Actual Quantity-Standard Quantity])* Standard Price]
=(187100-189000)*6    (11,400) F
Total Direct Materials Cost Variance (Direct Materials Price Variance+Direct Materials Quantity Variance)    (30,110) F
Direct Labor Rate Variance [(Actual rate per hour-Standard rate per hour)*Actual hours]
=(16.90-16.50)*17900       7,160 U
Direct Time Variance Variance [(Actual direct labor hours-Standard direct labor hours])* Standard rate per hour]
=(17900-17500)*16.5       6,600 U
Total Direct Labor Cost Variance (Direct Labor Rate Variance+Direct Time Variance Variance)     13,760 U
Variable factory overhead controllable variance Actual Variable OH    58,910
Budgeted Variable OH =(70,000*0.25)*3.40    59,500
Variable factory overhead controllable variance =         (590) F
Fixed factory overhead volume variance Budgeted Fixed OH    98,604
Fixed OH Cost applied (70000*0.25*5.40)    94,500
Fixed factory overhead volume variance =       4,104 U
Total factory overhead cost variance Variable Factory Overhead Controllable Variance + Fixed Factory Overhead Volume Variance       3,514 U